OpenStudy (anonymous):

Convert to an improper faction: 12 1/7 (Points : 1)

OpenStudy (anonymous):

Well to do this, you need the 12 to be into 7ths (common denominator). To do that you have to multiply 12 by 7. 12*7=84, and put this number as the numerator 84/7 = 12. (84/7)+(1/7)= (85/7)